# Learning Generalizable fMRI Representations for Neurological Disorder Identification

### v1

@@ 5.2 Evaluation of Representation Generalizability
We compare our representation learning method with foundation models to assess how well our approach generalizes to unseen clinical data relative to these powerful, large-scale models. Foundation models are designed to capture broad, generalizable representations across a wide variety of downstream tasks. By contrasting MeTSK’s learned representations with those from foundation models, we can evaluate the effectiveness of our method in generating robust and generalizable features that perform well even with limited clinical data.
We compared our MeTSK model to a large pre-trained fMRI model, as detailed in (Thomas et al., 2022). This model involves pre-training a Generative Pretrained Transformer (GPT) (Radford et al., 2019) on extensive datasets comprising 11,980 fMRI runs from 1,726 individuals across 34 datasets. During pre-training, the GPT model performs a self-supervised task to predict the next masked time point in the fMRI time-series. Their pre-trained model is publicly available at https://github.com/athms/learning-from-brains. We directly applied their pre-trained model to generate features. BrainLM (Ortega Caro et al., 2023), one of the latest foundation models developed for fMRI, was also incorporated into our experiments for comparison. BrainLM, consisting of a masked auto-encoder and a vision transformer, was trained on 6,700 hours of fMRI recordings. During pre-training, BrainLM incorporates a self-supervised task that predicts the randomly masked segments of time series in fMRI data, which is similar to the pre-training task in (Thomas et al., 2022). Similarly, we directly applied their pre-trained model available at https://github.com/vandijklab/BrainLM to generate features. We also pre-trained a ST-GCN model on both HCP and ADHD-Peking datasets using only the proposed contrastive self-supervised learning (SSL). From this pre-trained SSL model, we again directly generated features for unseen clinical data.
@@ 5.2.2 Evaluation Results
We compare the linear probing performance across foundation model approaches as well as the performance of classifiers trained with functional connectivity features extracted from raw fMRI data. We employed the same machine learning classifiers as used in the previous experiments, including a linear SVM, RF, and MLP. The same 5-fold cross-validation was applied and AUCs for classification tasks were computed. Although complex classifiers could be used, the use of SVM, RF and MLP isolates the quality of the learned representations from the model’s complexity.
Our proposed MeTSK achieved the best performance on all four datasets, outperforming the two latest fMRI foundation models, as shown in Table 2. The features generated by MeTSK, the SSL model and BrainLM (Ortega Caro et al., 2023) all achieved better performance than functional connectivity features, owing to the knowledge learned from their extensive pre-training datasets. However, due to the domain gap between their pre-training data and the clinical datasets used here, the foundation models may require additional fine-tuning to further improve their performance. Notably, MeTSK achieved statistically significant improvements in every dataset according to the p-values calculated using a paired t-test. These results highlight MeTSK’s potential in enhancing representation learning for clinical diagnostic purposes.
@@ 5.2.3 Interpretation of Learned Representations
Table 2: Generalization evaluation results using 5-fold cross-validation: Mean and std of AUCs for PTE, AD, and PD classification using representations generated from different models (denoted as "linear probing" in the Table) as well as directly from functional connectivity features. The best performance achieved on each dataset is highlighted in bold. Asterisks () indicate that the difference between our model MeTSK and the comparing methods is statistically significant using a paired Student’s t-test. Significance levels are denoted by ${{}^{*}(p\leq 0.05)}$, ${{}^{**}(p\leq 0.01)}$, ${{}^{***}(p\leq 0.001)}$.
@@
PTE SVM $\mathbf{0.6415\pm 0.0312}$ $0.5972\pm 0.0492^{*}$ $0.5369\pm 0.0451^{**}$ $0.6011\pm 0.0465$ $0.5697\pm 0.0477^{*}$
RF $0.5392\pm 0.0553$ $0.5253\pm 0.0486$ $0.4814\pm 0.0664^{*}$ $0.5589\pm 0.0611$ $0.5081\pm 0.0612$
MLP $0.5813\pm 0.0504$ $0.5216\pm 0.0329^{*}$ $0.5278\pm 0.0643$ $0.5290\pm 0.0674$ $0.5111\pm 0.0402^{*}$
OASIS-3 SVM $0.6407\pm 0.0741$ $0.5992\pm 0.0867^{*}$ $0.6115\pm 0.0582$ $0.6028\pm 0.0831$ $0.5541\pm 0.0677^{**}$
RF $\mathbf{0.6750\pm 0.0753}$ $0.6055\pm 0.0682^{*}$ $0.6233\pm 0.0672$ $0.5604\pm 0.0607^{**}$ $0.5329\pm 0.0721^{**}$
MLP $0.6034\pm 0.0725$ $0.5113\pm 0.0591^{**}$ $0.5593\pm 0.0576^{*}$ $0.5523\pm 0.0711^{**}$ $0.5231\pm 0.0689^{**}$
TaoWu SVM $\mathbf{0.6831\pm 0.1431}$ $0.6371\pm 0.1578$ $0.5528\pm 0.1586^{***}$ $0.4937\pm 0.1506^{***}$ $0.5725\pm 0.1502^{**}$
RF $0.6553\pm 0.1701$ $0.6273\pm 0.1679$ $0.4843\pm 0.1885^{***}$ $0.4891\pm 0.1710^{***}$ $0.6031\pm 0.1631$
MLP $0.6208\pm 0.1582$ $0.5013\pm 0.1621^{**}$ $0.5325\pm 0.1672^{*}$ $0.5078\pm 0.1592^{**}$ $0.5875\pm 0.1631$
Neurocon SVM $\mathbf{0.7529\pm 0.1579}$ $0.5643\pm 0.1068^{***}$ $0.6433\pm 0.1535^{***}$ $0.6476\pm 0.1686^{**}$ $0.6599\pm 0.1807^{*}$
RF $0.6230\pm 0.1658$ $0.5219\pm 0.1627^{***}$ $0.5813\pm 0.1818$ $0.6774\pm 0.1676$ $0.5427\pm 0.1715^{*}$
MLP $0.6100\pm 0.1635$ $0.5073\pm 0.1516^{**}$ $0.5401\pm 0.1802^{*}$ $0.5744\pm 0.1629$ $0.5313\pm 0.1721^{*}$
@@
Figure 4: A comparison of the linear probing performance between MeTSK model trained with a single target site (MeTSK-ADHD-Peking, results from Table 2) and MeTSK model trained using all the target sites from ADHD-200 dataset (MeTSK-ADHD-All). Also shown are results from linear classifiers trained directly using connectivity features for baseline comparison. The height of each bar indicates the average AUC computed from a 5-fold cross-validation, while the error bars denote the standard deviations. MeTSK-ADHD-All achieved similar performance to MeTSK-ADHD-Peking on all the four downstream datasets. The p-values computed using a paired Student’s t-test indicate that there is no significant difference in performance.
@@ 6.1 Training with More Clinical Data
However, while MeTSK is intentionally tailored for training with small-scale clinical data, we also explored the impact of using more clinical target data to train the model. We used the data from all of the sites in the ADHD-200 datasets for training. The entire dataset comprised a total of 362 ADHD subjects and 585 TDCs, all pre-processed using the same steps as described in Section 4.1.2. The same linear probing was performed to evaluate the models on the four clinical datasets: PTE, OASIS-3, TaoWu, and Neurocon. As shown in Fig 4, MeTSK consistently demonstrated similar performance for various downstream tasks when trained using the larger clinical dataset. We computed the p-value using a paired Student’s t-test, and showed that there is no significant difference between the downstream performance of the model trained with Peking site only (MeTSK-ADHD-Peking) and the model trained with the entire ADHD-200 dataset (MeTSK-ADHD-All). This underscores the model’s robustness and its capacity to achieve effective generalization without relying on a large amount of training data.
@@ 6.2 Ablation Study of MeTSK
Table 3: Ablation study on ADHD-Peking and ABIDE-UM dataset. The FT, MTL, and MeTSK methods are compared for two cases - transferring features from (i) a self-supervised source task and (ii) a sex classification source task, respectively. The last two rows are models trained only on target clinical data: a meta-learning model without source task and a baseline model.
@@ 7 Discussion
By employing meta-learning, we enhance the generalization capabilities of self-supervised features from the source domain (control) to target clinical domains, ensuring that the learned representations capture intrinsic patterns in functional brain activity, such patterns are shared across datasets. The learned representations enable effective classification even with simple linear classifiers. This demonstrates the quality of the learned features and their generalizability across unseen clinical datasets.
@@ A.1 Applying Deep Learning Models to Small-Scale Clinical Datasets
We present the classification results on the clinical datasets mentioned in Section 5.2 using some popular deep learning models. These models demonstrated severe over-fitting and poor generalization, achieving lower AUCs than traditional machine learning classifiers. The deep learning models struggled to learn from the limited and highly variable clinical data, where the high-dimensional and complex nature of fMRI data and small sample sizes amplified the over-fitting issues. These results emphasized our motivation for developing a representation learning approach capable of generating generalizable features for such challenging datasets.
@@
A.2 Experiment Results Using Additional Target Datasets

### v2

@@ 5.2 Evaluation of Representation Generalizability
We compare our representation learning method with foundation models to assess how well our approach generalizes to unseen clinical data relative to these powerful, large-scale models. Foundation models are designed to capture generalizable representations across a wide variety of downstream tasks. By contrasting MeTSK’s learned representations with those from foundation models, we can evaluate the effectiveness of our method in generating robust and generalizable features that perform well even with limited clinical data.
We compared our MeTSK model to a large pre-trained fMRI model, as detailed in Thomas et al. (2022). This model involves pre-training a Generative Pretrained Transformer (GPT) (Radford et al., 2019) on extensive datasets comprising 11,980 fMRI runs from 1,726 individuals across 34 datasets. During pre-training, the GPT model performs a self-supervised task to predict the next masked time point in the fMRI time-series. Their pre-trained model is publicly available at https://github.com/athms/learning-from-brains. We directly applied their pre-trained model to generate features. BrainLM (Ortega Caro et al., 2023), one of the latest foundation models developed for fMRI, was also incorporated into our experiments for comparison. BrainLM, consisting of a masked auto-encoder and a vision transformer, was trained on 6,700 hours of fMRI recordings. During pre-training, BrainLM incorporates a self-supervised task that predicts the randomly masked segments of time series in fMRI data, which is similar to the pre-training task in Thomas et al. (2022). Similarly, we directly applied their pre-trained model available at https://github.com/vandijklab/BrainLM to generate features. We also pre-trained a ST-GCN model on both HCP and ADHD-Peking datasets using only the proposed contrastive self-supervised learning (SSL). From this pre-trained SSL model, we again directly generated features for unseen clinical data.
@@ 5.2.2 Evaluation Results
We compare the linear probing performance across foundation model approaches as well as the performance of classifiers trained with functional connectivity features extracted from raw fMRI data. We employed machine learning classifiers, including a linear SVM, RF, and Logistic Regression (LR). The same 5-fold cross-validation was applied and AUCs for classification tasks were computed. Although complex classifiers could be used, the use of SVM, RF and LR isolates the quality of the learned representations from the model’s complexity.
Our proposed MeTSK achieved the best performance on all four datasets, outperforming the two latest fMRI foundation models, as shown in Table 2. The features generated by MeTSK, the SSL model and BrainLM (Ortega Caro et al., 2023) all achieved better performance than functional connectivity features, owing to the knowledge learned from their extensive pre-training datasets. However, due to the domain gap between their pre-training data and the clinical datasets used here, the foundation models may require additional fine-tuning to further improve their performance. These results highlight MeTSK’s potential in enhancing representation learning for clinical diagnostic purposes.
@@ 5.2.3 Interpretation of Learned Representations
Table 2: Generalization evaluation results using 5-fold cross-validation: Mean and std of AUCs for PTE, AD, and PD classification using representations generated from different models (denoted as "linear probing" in the Table) as well as directly from functional connectivity features. The best performance achieved on each dataset is highlighted in bold.
@@
PTE SVM $\mathbf{0.6415\pm 0.0312}$ $0.5972\pm 0.0492$ $0.5369\pm 0.0451$ $0.6011\pm 0.0465$ $0.5697\pm 0.0477$
RF $0.5392\pm 0.0553$ $0.5253\pm 0.0486$ $0.4814\pm 0.0664$ $0.5589\pm 0.0611$ $0.5081\pm 0.0612$
LR $0.5770\pm 0.0635$ $0.5087\pm 0.0455$ $0.5199\pm 0.0678$ $0.5378\pm 0.0663$ $0.5190\pm 0.0566$
OASIS-3 SVM $0.6407\pm 0.0741$ $0.5992\pm 0.0867$ $0.6115\pm 0.0582$ $0.6028\pm 0.0831$ $0.5541\pm 0.0677$
RF $\mathbf{0.6750\pm 0.0753}$ $0.6055\pm 0.0682$ $0.6233\pm 0.0672$ $0.5604\pm 0.0607$ $0.5329\pm 0.0721$
LR $0.5894\pm 0.0544$ $0.5044\pm 0.0581$ $0.5543\pm 0.0439$ $0.5207\pm 0.0437$ $0.5482\pm 0.0520$
Taowu SVM $\mathbf{0.6831\pm 0.1431}$ $0.6371\pm 0.1578$ $0.5528\pm 0.1586$ $0.4937\pm 0.1506$ $0.5725\pm 0.1502$
RF $0.6553\pm 0.1701$ $0.6273\pm 0.1679$ $0.4843\pm 0.1885$ $0.4891\pm 0.1710$ $0.6031\pm 0.1631$
LR $0.5725\pm 0.1696$ $0.5100\pm 0.1674$ $0.5175\pm 0.1848$ $0.5387\pm 0.1985$ $0.5275\pm 0.2057$
Neurocon SVM $\mathbf{0.7529\pm 0.1579}$ $0.5643\pm 0.1068$ $0.6433\pm 0.1535$ $0.6476\pm 0.1686$ $0.6599\pm 0.1807$
RF $0.6230\pm 0.1658$ $0.5219\pm 0.1627$ $0.5813\pm 0.1818$ $0.6774\pm 0.1676$ $0.5427\pm 0.1715$
LR $0.5885\pm 0.1815$ $0.5033\pm 0.1768$ $0.5238\pm 0.2001$ $0.5277\pm 0.1911$ $0.5163\pm 0.2167$
@@
Figure 4: A comparison of the linear probing performance between MeTSK model trained with a single target site (MeTSK-ADHD-Peking, results from Table 2) and MeTSK model trained using all the target sites from ADHD-200 dataset (MeTSK-ADHD-All). Also shown are results from linear classifiers trained directly using connectivity features for baseline comparison. The height of each bar indicates the average AUC computed from a 5-fold cross-validation, while the error bars denote the standard deviations. MeTSK-ADHD-All achieved similar performance to MeTSK-ADHD-Peking on all the four downstream datasets.
5.3 Implementation Details
Training: To optimize model performance, we follow the training setting in (Gadgil et al., 2020) for the ST-GCN model. The length of input sub-sequences for ST-GCN is fixed at 128. We generate one meta-training batch by randomly selecting an equal number of samples from each class. The batch size is 32, both for the meta-training and the meta-validation set. We use an Adam optimizer (Kingma & Ba, 2014) with learning rate $\beta=0.001$ in the outer loop, and an SGD optimizer (Ketkar, 2017) with learning rate $\alpha=0.01$ in the inner loop. The number of inner loop update steps is $25$. We set the hyper-parameter $\lambda=30$ and the temperature parameter $\tau=30$ to adjust the scale of losses following (Liu et al., 2020a; You et al., 2020a). Since contrastive loss converges slowly (Jaiswal et al., 2020), a warm-up phase is applied to train the model only on HCP data using the graph contrastive loss for the first 20 epochs. The number of total training epochs is 90. The hyper-parameters for the baseline ST-GCN model and the contrastive SSL loss follow the settings reported in Gadgil et al. (2020); You et al. (2020a). Other hyper-parameters in MeTSK were selected empirically by observing the convergence of the training loss. The hyper-parameters in SVM and RF during linear probing are selected via grid search with nested cross-validation. For the multi-task learning (MTL) implementation, we simply remove the inner loop in MeTSK and use all the training data to update the target head. Both heads and the feature extractor are updated simultaneously in one loop. For meta-learning, the target training set in each fold is further divided into a meta-training set $X_{\mathcal{T}_{tr}}$ of $157$ subjects and a meta-validation set $X_{\mathcal{T}_{val}}$ of $39$ subjects. Both the meta-train and meta-validation sets are randomly re-sampled in every iteration, and the target head is randomly initialized. In the training of MeTSK, binary classification is performed using one fully-connected layer followed by a sigmoid activation. The classification loss is the binary cross-entropy loss. .
Evaluation: For SSL and MeTSK, we use the pre-trained feature extractor for generating features. The generated features are graph-level representations, having a two dimensional feature matrix at each node (brain region). We averaged the features along the first dimension and applied Pinciple Component Analysis (PCA) to reduce the dimensionality before feeding the features into classifiers. The logistic regression used in the downstream experiments consists of a linear layer and followed by sigmoid activation. The SSL model trained on both HCP and ADHD-Peking data used the same contrastive loss and hyper-parameters in MeTSK. In our comparative analysis with a foundation model for fMRI (Thomas et al., 2022), we flatten the brain signals at each time-point and input the whole time-series without masking into the pre-trained GPT model. This generates a feature embedding for each time-point. Since there is no class token in the pre-trained model, we averaged the feature embeddings across time-points (tokens) to get the output features for downstream datasets. We follow the other detailed settings of the pre-trained GPT model in (Thomas et al., 2022). For another foundation model BrainLM (Ortega Caro et al., 2023), we followed the instructions provided on https://github.com/vandijklab/BrainLM/tree/main. Specifically, we directly input the fMRI time-series into the pre-trained model without masking. Then we extracted the output class token as the generated features for clinical datasets.
Computational Cost: MeTSK introduces additional training time because there is a nested optimization loop within each training iteration. Training MeTSK takes approximately 6 hours on a single NVIDIA V100 GPU with 32 GB of memory. This nested gradient computation in the inner loop increases both training time and memory consumption. Larger GPU memory is recommended when scaling MeTSK to larger datasets or models. The SSL model requires about 4 hours of training using the same GPU. Baseline models such as ST-GCN, LSTM, and STAGIN typically take between 1.5 to 2 hours to train. Since we evaluate the two foundation models using pre-trained checkpoints, their inference cost is negligible. While MeTSK has higher training cost, it is a one-time cost at the pretraining stage. Once trained, the representations can be reused across downstream tasks with minimal computational overhead using linear probing.
@@ 6.1 Training with More Clinical Data
However, while MeTSK is intentionally tailored for training with small-scale clinical data, we also explored the impact of using more clinical target data to train the model. We used the data from all of the sites in the ADHD-200 datasets for training. The entire dataset comprised a total of 362 ADHD subjects and 585 TDCs, all pre-processed using the same steps as described in Section 4.1.2. The same linear probing was performed to evaluate the models on the four clinical datasets: PTE, OASIS-3, TaoWu, and Neurocon. As shown in Fig 4, MeTSK consistently demonstrated similar performance for various downstream tasks when trained using the larger clinical dataset. We computed the p-value using a paired Student’s t-test, and showed that there is no significant difference between the downstream performance of the model trained with Peking site only (MeTSK-ADHD-Peking) and the model trained with the entire ADHD-200 dataset (MeTSK-ADHD-All). This underscores the model’s robustness and its capacity to achieve effective generalization without relying on a large amount of training data. In some cases, training with the entire ADHD dataset achieved worse performance. The ADHD-All dataset includes fMRI scans from multiple imaging sites with different acquisition protocols, which can introduce distribution shifts. This phenomenon has been studied in prior work (the data addition dilemma) Shen et al. (2024), which highlights that adding more data from heterogeneous sources can in some cases degrade model performance due to domain shifts.
@@ 6.2 Ablation Study of MeTSK
Table 3: Ablation study on ADHD-Peking and ABIDE-UM dataset. The FT, MTL, and MeTSK methods are compared for two cases - transferring features from (i) a self-supervised source task and (ii) a sex classification source task, respectively. The last two rows are models trained only on target clinical data: a meta-learning model without source task and a baseline model ST-GCN.
@@ 7 Discussion
In this work, we primarily focus on evaluating representation generalization in low-data regimes. Although some datasets, such as OASIS-3, contain a large number of subjects, we only used a subset of the entire dataset and applied a 5-fold cross-validation protocol to maintain consistency across all datasets. Cross-validation provides a balanced estimate of generalization performance across multiple folds and ensures the evaluation setup is consistent across datasets, including those with naturally limited sample sizes. Although training on a small subset and evaluating on the entire dataset is interesting to explore, it remains challenging because the limited training subjects may not capture the full variability of the entire population.
Further explorations exist for expanding this approach in future work. One direction is to experiment with alternative functional brain atlases, such as DiFuMO (Dadi et al., 2020), or to incorporate different functional connectivity measures (e.g. Partial correlation) beyond Pearson’s correlation (Varoquaux & Craddock, 2013; Smith et al., 2011). Note that our current approach is compatible with various choices of atlases and connectivity measures, which opens the potential to extend MeTSK to more advanced connectivity metrics and to easily adapt it to different atlas configurations. Additionally, exploring multi-class classification for distinguishing between multiple neurological disorders in downstream tasks rather than a simple binary classification could potentially provide further insights.
8 Conclusion
To summarize, we developed a representation learning strategy for neurological disorders, which achieved superior generalization to diverse small-scale clinical datasets and significantly improved the performance on various challenging neurological disorder classification tasks. By employing meta-learning, we enhance the generalization capabilities of self-supervised features from the source domain (control) to target clinical domains, ensuring that the learned representations capture intrinsic patterns in functional brain activity and such patterns are shared across datasets. The learned representations enable effective classification even with simple linear classifiers. This demonstrates the quality of the learned features and their generalizability across unseen clinical datasets.
@@ A.1 Applying Deep Learning Models to Small-Scale Clinical Datasets
We present the classification results on the clinical datasets mentioned in Section 5.2 using some popular deep learning models. These models demonstrated severe over-fitting and poor generalization, achieving lower AUCs than traditional machine learning classifiers. The deep learning models struggled to learn from the limited and highly variable clinical data, where the high-dimensional and complex nature of fMRI data and small sample sizes amplified the over-fitting issues. These results shown in Table 4 emphasized our motivation for developing a representation learning approach capable of generating generalizable features for such challenging datasets.
@@
Due to methodological and experiment setting differences across previous works (e.g., varying cross-validation splits, fMRI atlas choices, and number of training samples used), direct numerical comparisons should be interpreted cautiously. For example, different fMRI atlases define varying numbers of ROIs, which directly impact the dimensionality of the input connectivity features and change the number of nodes in the input graphs. We include these published results from previous works to contextualize our results within the broader literature, as shown in Table 5. Unlike previous works that often leverage larger datasets, our experimental design simulates real-world clinical scenarios with limited training data. This allows us to assess the robustness and generalization capability of MeTSK under data-scarce conditions—settings that are common in clinical practice.
Table 5: Summary of reported results on our four clinical datasets in the literature as well as the best results achieved by MeTSK. Note the significant differences in metrics, cross-validation splits, training size, and fMRI atlases.
Dataset Method Results Metric CV Splits # Training Samples fMRI Atlas # ROIs
Neurocon Xu et al. (2024) 0.75 Accuracy 10-fold CV 42 Schaefer (Schaefer et al., 2018) 100
Shi et al. (2022) 0.667 AUC 10-fold CV 42 Brainnetome (Fan et al., 2016) 246
Vigneshwaran et al. (2024) 0.5 F1 Score 40% for testing 42 Harvard-Oxford 69
Li et al. (2021) 0.66 Accuracy 10-fold CV 42 Schaefer 100
MeTSK (Ours) 0.7529 AUC 5-fold CV 42 AAL 116
Taowu Xu et al. (2024) 0.775 Accuracy 10-fold CV 40 Schaefer 100
Vigneshwaran et al. (2024) 0.6 F1 Score 40% for testing 40 Harvard-Oxford Atlas (Wu et al., 2018) 69
Li et al. (2021) 0.675 Accuracy 10-fold CV 40 Schaefer 100
MeTSK (Ours) 0.6831 AUC 5-fold CV 40 AAL 116
OASIS-3 Han et al. (2024) (Proposed GCN) 0.800 Accuracy 10-fold CV 1006 Schaefer 100
Han et al. (2024) (Baseline GCN) 0.686 Accuracy 10-fold CV 1006 Schaefer 100
Han et al. (2024) (SVM) 0.568 Accuracy 10-fold CV 1006 Schaefer 100
MeTSK (Ours) 0.675 AUC 5-fold CV 84 AAL 116
PTE Akrami et al. (2024) (KSVM) 0.6100 AUC Leave-one-out CV 72 USCLobes (Joshi et al., 2022) 16
Akrami et al. (2024) (GCN) 0.5900 AUC Leave-one-out CV 72 USCLobes 16
MeTSK (Ours) 0.6415 AUC 5-fold CV 72 AAL 116
A.2 P-values Computed from Wilconxon Signed-Rank Tests
We report the p-values computed using the Wilcoxon signed-rank test between the results from Connectivity Features, MeTSK-ADHD-All, and MeTSK-ADHD-Peking in Table 6. As shown, the comparisons between MeTSK-ADHD-Peking and MeTSK-ADHD-All consistently yield high p-values, indicating no statistically significant difference between them. In contrast, comparisons between MeTSK-ADHD-Peking and Connectivity Features result in smaller p-values, suggesting more substantial differences in performance.
Table 6: Wilcoxon signed-rank test p-values between MeTSK and two other methods in Figure 4 across all datasets and classifiers. Columns represent comparisons against MeTSK-ADHD-Peking using MeTSK-ADHD-All and Connectivity Features, respectively. A higher p-value indicates a less significant difference between methods.
Dataset Classifier vs ADHD-All vs Connectivity Features
PTE SVM 0.4375 0.3125
RF 0.8125 0.4375
LR 1.0 0.3125
OASIS-3 SVM 0.625 0.1250
RF 0.8125 0.1250
LR 0.8125 0.3125
TaoWu SVM 0.8125 0.1250
RF 0.8125 0.3125
LR 0.625 0.3125
Neurocon SVM 0.8125 0.3125
RF 0.4375 0.3125
LR 0.675 0.3125
A.3 Convergence Analysis of MeTSK
To monitor the convergence and stability of MeTSK training, we track both the self-supervised loss and the cross-entropy loss computed from the frozen target head’s predictions using the meta-validation set in Step 3 (refer to Method Section 3.2). We plot the evolution of the two losses over training iterations in Figure 5. During training, we observe that the self-supervised loss decreases well, while the meta-validation cross-entropy (classification) loss also shows a decreasing trend overall, although with some fluctuations. Fluctuations are expected, as the loss is computed on a small (39-subject) randomly sampled meta-validation set. Both the meta-train and meta-validation sets are re-sampled in each training iteration, and the target classification head is randomly initialized, introducing further variability in the observed loss values. This indicates that the model is able to jointly optimize the self-supervised and classification losses in Step 3, and that the convergence is not negatively affected by the frozen target head trained in Step 2.
The classification loss on the meta-training set after inner-loop updates in Step 2 consistently decreases to a stable range across training epochs, as shown in Figure 5. Additionally, the final meta-training classification accuracies after inner-loop updates fall within a stable range of 72% to 81%. This suggests that the target head can quickly adapt to the updated features and the optimization steps in Step 3 do not negatively affect the convergence of the inner-loop updates in Step 2.
Figure 5: Loss evolution during MeTSK training over 90 epochs. The first 20 epochs is a warm-up phase where only the self-supervised task is trained. All loss values are averaged over batches and recorded every 2 epochs. Left: Classification loss on the meta-training set after the inner-loop updates in Step 2. We record the loss at the last update step in each training iteration. This loss remains within a stable range and gradually decreases, indicating that the target head effectively adapts using features trained in Step 3 without negative interference. Middle: Self-supervised loss from Step 3, showing a steady decline. Right: Meta-validation classification loss in Step 3, computed from the frozen target head, exhibits a generally decreasing trend with some fluctuations.
A.4 Linear Probing Results of Fine-tuned Foundation Models
To ensure a fair comparison, we fine-tuned the baseline foundation models on the same dataset—HCP combined with ADHD-Peking—that was used for training MeTSK. Notably, the pre-training datasets for these foundation models already included the HCP data, so aligning the training data required only fine-tuning on the ADHD-Peking dataset. Specifically, we load the pre-trained foundation models and fine-tune them on the ADHD-Peking dataset using the same self-supervised pre-training objective originally used to train the models—reconstructing masked time points. After fine-tuning on ADHD-Peking, we performed linear probing on the four clinical datasets. Our results in Table 7 show that the linear probing performance of the fine-tuned foundation models is similar to or slightly worse than without finetuning. This may be due to the large number of parameters in the foundation models (e.g., BrainLM has 13 million parameters), which makes them more prone to overfitting when fine-tuned on the relatively small ADHD-Peking dataset (245 subjects versus over 40,000 in the pre-training datasets of baseline foundation models).
Table 7: Comparison of linear probing performance from foundation models Thomas et al. (2022); Ortega Caro et al. (2023) and their fine-tuned versions on four clinical datasets using 5-fold cross-validation. We directly used results of the foundation model w/o finetuning from Table 2 for comparison. Mean $\pm$ std AUCs are reported for both pre-trained and fine-tuned models. Fine-tuned results are obtained by further training the foundation models on the ADHD-Peking dataset.
Dataset Classifier Thomas et al. (2022) Fine-tuned Ortega Caro et al. (2023) Fine-tuned
PTE SVM $0.5369\pm 0.0451$ $0.5291\pm 0.0473$ $0.6011\pm 0.0465$ $0.6102\pm 0.0498$
RF $0.4814\pm 0.0664$ $0.4795\pm 0.0651$ $0.5589\pm 0.0611$ $0.5502\pm 0.0584$
LR $0.5199\pm 0.0678$ $0.5252\pm 0.0662$ $0.5378\pm 0.0663$ $0.5415\pm 0.0691$
OASIS-3 SVM $0.6115\pm 0.0582$ $0.6152\pm 0.0603$ $0.6028\pm 0.0831$ $0.5955\pm 0.0792$
RF $0.6233\pm 0.0672$ $0.6184\pm 0.0660$ $0.5604\pm 0.0607$ $0.5601\pm 0.0622$
LR $0.5543\pm 0.0439$ $0.5482\pm 0.0455$ $0.5207\pm 0.0437$ $0.5280\pm 0.0411$
TaoWu SVM $0.5528\pm 0.1586$ $0.5477\pm 0.1602$ $0.4937\pm 0.1596$ $0.4913\pm 0.1530$
RF $0.4843\pm 0.1885$ $0.4782\pm 0.1903$ $0.4891\pm 0.1710$ $0.4965\pm 0.1684$
LR $0.5175\pm 0.1848$ $0.5083\pm 0.1822$ $0.5387\pm 0.1985$ $0.5280\pm 0.1971$
Neurocon SVM $0.6433\pm 0.1535$ $0.6382\pm 0.1510$ $0.6476\pm 0.1686$ $0.6488\pm 0.1755$
RF $0.5813\pm 0.1818$ $0.5755\pm 0.1799$ $0.6774\pm 0.1676$ $0.6703\pm 0.1728$
LR $0.5238\pm 0.2001$ $0.5104\pm 0.2153$ $0.5277\pm 0.1911$ $0.5292\pm 0.1984$
A.5 Experiment Results Using Additional Target Datasets